Suppose real Gdp is growing at 4%, the money supply is growing at 11%, the velocity of money is constant,and the interest rate is 6%.
a) Whats the current inflation rate and norminal interest rate?
b) If the money supply growth rate increases to 15%, how will your answer in part (a) change?
c) If you were an investor, how would the change in the money supply growth affect your real profitability, assuming that you now receive the new norminal interest rate?

d) Based on previous answers, would you prefer a fixed or floating interest rate on your investments? which would you prefer if you thought the money supply growth was going to be reduced?

ANSWER FOR ME (c) and (d) please
1
Expert's answer
2012-09-07T11:29:04-0400
i) M/P=Y/V
Y - growing by 4% so coeficient = 1.04
M - growing by 11% so coeficient = 1.11
V - constant =1
1.11/P=1.04/1
P=1.11/1.04≈1.067≈+6.7%

r-real interest rate
i-nominal interest rate
π-inflation
(1+r)=(1+i)/(1+π)
1.06=(1+i)/(1+0.067)
i=(1.06*1.067)-1
i≈0.13102≈13.10%
______ ______ ______ ______

(ii)
M/P=Y/V
Y - growing by 4% so coeficient = 1.04
M - growing by 15% so coeficient = 1.15
V - constant =1
1.15/P=1.04/1
P=1.15/1.04≈1.1057692≈+10.57%

r-real interest rate
i-nominal interest rate
π-inflation
(1+r)=(1+i)/(1+π)
1.06=(1+i)/(1+0.1057)
i=(1.06*1.1057)-1
i≈0.1720≈17.20%
______ ______ ______ ______

i¹≈13.10%
i²≈17.20%
Δi=(17.20/13.10)-1≈
≈1.3129-1≈+31.29%